# Graph sub-sampling for divide-and-conquer algorithms in large networks

### v3

@@ 5.1 Community detection
[…]
We apply the $\mathsf{PACE}$ algorithm to each network using Fast Greedy (Clauset et al.,, 2004) as the base community detection algorithm. We report the proportion of nodes in the largest community, $\hat{\alpha}$, as well as the modularity of the returned node membership vector, $Q$ (Newman,, 2006). The modularity is a measure of the strength of community structure in a network where a large value indicates a better community assignment. For a network ${\bf A}$ and community labels $\boldsymbol{c}$, it is computed as
@@
where $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$. Since we have the ground truth labels (which we call the Oracle) for the Political Blogs network, we can also compute the ARI between the true and estimated labels. Finally, we set $q=250/n$ for each network, and $B=1000$ and $5000$ for Political Blogs and Facebook, respectively.
@@ 5.2 Core-periphery detection
[…]
@@
[…]
In Table 4, we report the core size and BE metric for the detected cores, and in Figure 5, we compare the cores returned by the different algorithms. Specifically, for each sub-sampling pair, we compute the Jacaard coefficient (JC) between the returned core sets. If $S_{1}$ and $S_{2}$ are two sets, then the JC is computed as
@@ 6 Conclusion
It is important to note that these results were based on our simulated and real-world data experiments and may not generalize to other settings. That being said, the contrasting performance of sub-sampling routines on the different identification task is revealing. Random node and random node neighbor sampling performed the best for the community detection problem, but were the two weakest methods for core-periphery identification. On the other hand, breadth-first and depth-first searches are good approaches for identifying core-periphery structure, but not for community structure. These results highlight the importance of tailoring the choice of sub-sampling algorithm to the particular problem. Indeed, our CP theory revealed the optimal feature for a sampler is a high probability of sampling core nodes. Deriving similar theory for future problems will help researchers choose the best methods, or even may even drive them to derive novel sub-sampling schemes for the specific problem.
@@
$\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{n}(\mathbb{E}||\hat{\boldsymbol{c}}^{(S)}-\boldsymbol{c}^{*(S)}||_{2}^{2}+\mathbb{E}\sum_{i=1}^{n}y_{i}^{(S)})$ (10)
@@
Corollary 3.2:
Let ${\bf A}$ be generated from a CP-SBM with $n$ nodes and $k$ core nodes. Let $S$ be a randomly chosen sub-graph of ${\bf A}$ of size $q_{n}n$ where $q_{n}\in(0,1)$ may depend on $n$. Furthermore, let $y_{i}^{(S)}$ is the event that node $i$ is a core node and not sampled in $S$. If RN is used to sample $S$ and $q_{n}=O(1)$, then
@@
$\lim_{n\to\infty}\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}=\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n}.$ (12)
@@
Random Node: Recall that $q_{n}=q=O(1)$ in this setting. Let $Y^{(\mathcal{S})}$ be the number of core nodes drawn in sub-graph $\mathcal{S}$, i.e.,
@@
Degree Node: When we sample the first node, the probability that it is a core node is
@@
Random Edge: We follow a similar pattern for RE. Recall that we are sampling edges now, so the probability that the first edge we sample is incident to a core node is:
@@
We replicate the simulation study from the main manuscript, now using the Walktrap community detection algorithm (Pons and Latapy,, 2005) as the base algorithm. We use the same settings as in the main manuscript. The results are in Figures 6 and 7. We can see that the results are very similar to those with the Fast Greedy base algorithm. The main difference is that the base algorithm outperforms $\mathsf{PACE}$ for all settings (except extremely disproportionate community sizes), but is slower for many settings. These results are more in line with what we would expect: the divide-and-conquer algorithm yields slightly worse identification results but is computationally faster.
@@
In Figures 8 and 9, we report the average time it takes to detect the community structure and CP structure on the sub-samples, respectively, for the main simulation settings. In general, RNN is the slowest with RN being the fastest.
@@
where ${{\bf A}}^{(b)}$ is the adjacency matrix for sub-sampled network $\mathcal{S}^{(b)}$. Note that settings 3 (9) and 5 (11) are redundant for the community (CP) simulations so they are excluded. We also plot the density of the original graph. The results are in Figures 10 and 11.

### v4

@@ 5.1 Community detection
We apply the $\mathsf{PACE}$ algorithm to two different real-world networks. This first is Political Blogs (Adamic and Glance,, 2005) ($n=1\hskip 1.9919pt490,m=16\hskip 1.9919pt715$) where nodes represent blogs and edges arise if one blog references another. The blogs are divided between liberal and conservative, so we set $K=2$, and keep only the largest connected component, yielding $n=1222$. This network was chosen since there are ground-truth labels (conservative or liberal) that we can compare with the estimated labels, and also because it was considered in the original $\mathsf{PACE}$ paper. The second is Facebook (Leskovec and Mcauley,, 2012) ($n=4\hskip 1.9919pt039,m=88\hskip 1.9919pt234$) where nodes are users connected with an edge if they are friends. This is an ego network with ten egos, so we set $K=10$. An ego refers to the sub-graph of all friends of a single user. This network was chosen since the number of communities is known, a requirement for $\mathsf{PACE}$. We remove all edge weights and directions as well as self-loops for both networks.
We apply the $\mathsf{PACE}$ algorithm to each network using Fast Greedy (Clauset et al.,, 2004) as the base community detection algorithm. We report the proportion of nodes in the largest community, $\hat{\alpha}$, as well as the modularity of the returned node membership vector, $Q$ (Newman,, 2006). Modularity is a popular measure of the strength of community structure where a large value indicates a better community assignment (e.g., Newman and Girvan,, 2004; Bickel and Chen,, 2009). For a network ${\bf A}$ and community labels $\boldsymbol{c}$, it is computed as
@@
where $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$. Since we have the ground truth labels (which we call the Oracle) for the Political Blogs network, we can also compute the ARI between the true and estimated labels. Finally, we set $q=250/n$ for each network, and $B=1\hskip 1.9919pt000$ and $5\hskip 1.9919pt000$ for Political Blogs and Facebook, respectively.
@@ 5.2 Core-periphery detection
We now compare the sub-sampling routines using the CP divide-and-conquer algorithm on real-world networks: Airport (Csárdi et al.,, 2025) is a network where nodes correspond to airports and edges to flights between airports ($n=755$, $m=4\hskip 1.9919pt623$); and Twitch (Rozemberczki et al.,, 2021; Leskovec and Krevl,, 2014) has nodes which are users and edges represent friendships ($n=168\hskip 1.9919pt114$, $m=6\hskip 1.9919pt797\hskip 1.9919pt557$). Edge weights and directions were again removed, along with self-loops.
@@
While the airport network is relatively small, this allows us to find the CP labels using the base algorithm without the divide-and-conquer step. Recall that the base algorithm is based on Algorithm 1 from Yanchenko, (2022). We label this as “Full” in our results and provides us with a useful benchmark. Finally, we note the following values of $q$ and $B$ that were used for each network: Airport, $q=100/n$ and $B=5\hskip 1.9919pt000$; and Twitch, $q=500/n$ and $B=25\hskip 1.9919pt000$.
In Table 4, we report the core size and BE metric for the detected cores, and in Figure 5, we compare the cores returned by the different algorithms. Specifically, for each sub-sampling pair, we compute the Jacaard coefficient (JC) (Liben-Nowell and Kleinberg,, 2003) between the returned core sets. If $S_{1}$ and $S_{2}$ are two sets, then the JC is computed as
@@ 6 Conclusion
It is important to note that these results were based on our simulated and real-world data experiments and may not generalize to other settings. That being said, the contrasting performance of sub-sampling routines on the different identification tasks is revealing. Random node and random node neighbor sampling performed the best for the community detection problem, but were the two weakest methods for core-periphery identification. On the other hand, breadth-first and depth-first searches are good approaches for identifying core-periphery structure, but not for community structure. These results highlight the importance of tailoring the choice of sub-sampling algorithm to the particular problem. Indeed, our CP theory revealed the optimal feature for a sampler is a high probability of sampling core nodes. Deriving similar theory for future problems will help researchers choose the best methods, or even may even drive them to derive novel sub-sampling schemes for the specific problem.
Acknowledgments
The author would like to thank the Associate Editor and two anonymous referees whose comments and suggestions greatly improved the manuscript.
@@ Acknowledgments
Sub-sampling algorithms
Figure 6 displays a diagram comparing the different sub-sampling algorithms. [^4] We are greatly indebted to one of the anonymous reviewers for providing this figure. In pane (a), we plot a small graph with $n=12$ nodes and $m=15$ edges. In (b) and (c), randomly sampled nodes are marked in orange with a black border, with their connecting edges also in black. Sampled edges are in orange while the incident nodes are colored in black with an orange border in (d). The first node of the search is represented with a white circle in the center for (e) and (f), with the other visited nodes colored in orange. The first node has the same white mark in (g) and (h). This figure reveals some insights into the similarities and differences in graph topology of the sub-graphs induced by the different sampling algorithms.
(a) Original graph
(b) Random node
(c) Degree node
(d) Random edge
(e) Breadth-first search
(f) Depth-first search
(g) Random node neighbor
(h) Random walk
Figure 6: Comparison of each of the seven sub-sampling algorithms.
@@
$\mathbb{E}\delta(\hat{\boldsymbol{c}},\boldsymbol{c}^{*})\leq\frac{1}{n}\left(\mathbb{E}||\hat{\boldsymbol{c}}^{(S)}-\boldsymbol{c}^{*(S)}||_{2}^{2}+\mathbb{E}\sum_{i=1}^{n}y_{i}^{(S)}\right)$ (10)
@@
$\mathbb{E}\frac{1}{n}\sum_{i=1}^{n}y_{i}^{(S)}\sim\left(1-q_{n}\frac{p_{12}}{p_{22}}\right)\alpha_{n},$ (12)
where $\sim$ denotes the asymptotic behavior for large $n$.
@@
(Random node) Recall that $q_{n}=q=O(1)$ in this setting. Let $Y^{(\mathcal{S})}$ be the number of core nodes drawn in sub-graph $\mathcal{S}$, i.e.,
@@
Proof. (Degree node) When we sample the first node, the probability that it is a core node is
@@
Proof. (Random edge) We follow a similar pattern for RE. Recall that we are sampling edges now, so the probability that the first edge we sample is incident to a core node is:
@@
We replicate the simulation study from the main manuscript, now using the Walktrap community detection algorithm (Pons and Latapy,, 2005) as the base algorithm. We use the same settings as in the main manuscript. The results are in Figures 7 and 8. We can see that the results are very similar to those with the Fast Greedy base algorithm. The main difference is that the base algorithm outperforms $\mathsf{PACE}$ for all settings (except extremely disproportionate community sizes), but is slower for many settings. These results are more in line with what we would expect: the divide-and-conquer algorithm yields slightly worse identification results but is computationally faster.
@@
In Figures 9 and 10, we report the average time it takes to detect the community structure and CP structure on the sub-samples, respectively, for the main simulation settings. In general, RNN is the slowest with RN being the fastest.
@@
where ${{\bf A}}^{(b)}$ is the adjacency matrix for sub-sampled network $\mathcal{S}^{(b)}$. Note that settings 3 (9) and 5 (11) are redundant for the community (CP) simulations so they are excluded. We also plot the density of the original graph. The results are in Figures 11 and 12.